King City Specialty Bikes (KCSB) produces high-end bicycles. The costs to manufacture and market the bicycles at the company's volume of 2,000 units per month are shown in the following table:

Unit manufacturing costs
Variable costs $ 250
Fixed overhead 117
Total unit manufacturing costs $ 367
Unit nonmanufacturing costs
Variable 70
Fixed 137
Total unit nonmanufacturing costs 207
Total unit costs $ 574

The company has the capacity to produce 2,000 units per month and always operates at full capacity. The bicycles sell for $610 per unit.

Required:

a. KCSB receives a proposal from an outside contractor who will assemble 800 of the 2,000 bicycles per month and ship them directly to KCSB’s customers as orders are received from KCSB’s sales force. KCSB would provide the materials for each bicycle, but the outside contractor would assemble, box, and ship the bicycles. The variable manufacturing costs would be reduced by 30 percent for the 800 bicycles assembled by the outside contractor. KCSB’s fixed nonmanufacturing costs would be unaffected, but its variable nonmanufacturing costs would be cut by 70 percent for these 800 units produced by the outside contractor. KCSB’s plant would operate at 60 percent of its normal level, and total fixed manufacturing costs would be cut by 25 percent.

a-1. Calculate the in-house unit cost that must be compared with the quotation received from the outside contractor. Assume the payment to the outside contractor is $150.

a-2. Should the proposal be accepted for a price (that is, payment to the contractor) of $150 per unit?

Yes
No

b. Assume the same facts as in requirement (a) but assume that the idle facilities would be used to produce 80 specialty racing bicycles per month. These racing bicycles could be sold for $7,700 each, while the costs of production would be $5,300 per unit variable manufacturing cost. Variable marketing costs would be $170 per unit. Fixed nonmanufacturing and manufacturing costs would be unchanged whether the original 2,000 regular bicycles were manufactured or the mix of 1,200 regular bicycles plus 80 racing bicycles was produced. What is the total net profit/loss for the following.

b-1. When the company produces and sells 2,000 units of regular bicycles per month. Assume the payment to the outside contractor is $150.

b-2. When the company produces 1,200 units of regular bicycles and use the idle facilities to produce 80 specially racing bicycles per month.
Answer #1
In-house cost to be compared is equal to the avoidable cost per unit
Unit Manufacturing cost
Variable 75
Fixed 73.125
Unit non manufacturing cost
Variable 49
Fixed 0
Avoidable unit cost 197.125
a-2. Yes, since the in-house cost is more
b-1Profit = (610-250-70)*1200 - 117*2000*75%-137*2000 + (610-150-175-21)*800
=$109,700
b-2 Profit = (610-250-70)*1200 - 117*2000-137*2000 + (610-150-175-21)*800 - (7700-5300-170)*80
=$229,600